# A Sparse Graph Formulation for Efficient Spectral Image Segmentation

### v1

@@ 4.2 Results and Discussion
In Table 1, we show the quantitative assessment of our proposed method for images with $\ell=100$ pixels and two different pairs of noise contaminations for generating each synthetic image under the three proposed ground-truth patterns. For our method, we depict the best JAC when considering the result from trying $\mu$ and $\sigma\in[0,1,\ldots,10]$. For the other methods, we tuned their available parameters and report their best JAC results. In the same table, we also display the average runtime for generating each of the best results’ performance in each method per pattern. In Figure 1, we qualitatively compare the best results of each method under two challenging noise levels. These results demonstrate that our method greatly over performs the comparative methods in terms of runtime, a consequence of its sparse nature, while also being competitive in terms of JAC, attaining almost perfect segmentations in most challenging scenarios.
In Figure 2, we qualitatively evaluate the effect of $\mu$ to the final segmentation. In these experiments, $\sigma$ is set to 0. In this same figure, we also qualitatively demonstrate our method’s performance and runtime in large RGB images. Here, we note that $\mu$ has the effective role of smoothing the borders of the final segmentation, removing artifacts or boundary details as it increases.
@@
Finally, in Figure 3 we present the average runtime and JAC values for the segmentations of the three patterns of them. We vary the image side lengths $\ell$ from $100$ to $1000$ and contaminated each image with zero mean Gaussian noise with $\sigma_{g}^{2}=\ell/100$, which means that images get increasingly hard with size. We again keep $\sigma=0$ in our proposed algorithm. These results demonstrate that our method scales well despite its spectral nature and generates high quality segmentations on large images even in very challenging noise settings.
5 Conclusion and Future Directions
@@
In this paper, we introduce a new spectral algorithm for image segmentation that overcomes some historical issues found in the Normalized Cuts framework and in its related methods. We employ an efficient sparse graph, whose usage under the traditional Normalized Cuts optimization problem leads to an interpretable cut value. We also show that our method is fast and produces high quality results for large RGB imagery and under high synthetic noise settings. For future work, we will derive an interpretation of the cut found on the graph proposed in Section 3.3.1 and extend our methodology to superpixel images—which can further still improve the overall scalability of our proposed method—and multi-region image segmentation.
[ scale only axis, height=1.5cm, width=.75ytick pos=left, ylabel=JAC, yticklabel style = font=, color=red, xticklabel style = font=, every y tick/.style=red, ylabel near ticks, xlabel near ticks, xmin=100,xmax=800, ymax=0.1, scaled y ticks=true, yticklabel style=/pgf/number format/fixed ]
[mark=none, red] table[x index=0,y index=2] data/size_times_average_variable.dat;
[…]
Fig. 3: Segmentation results for varying image sizes. Left: a sample image for $\ell=500$. Right: the runtime and JAC value results varying $\ell$ (on the $x$-axis).

### v3

@@ 4.2 Results and Discussion
In Table 1, we show the quantitative assessment of our proposed method for images with $\ell=100$ pixels and the two different noise contaminations when generating each synthetic image under the three proposed ground-truth patterns. We also display the average runtime for generating each of the best segmentations in each method per pattern. In Figure 1, we qualitatively compare the best results of each method under two challenging noise settings. These results demonstrate that our method greatly outperforms its related spectral methods in terms of runtime, a consequence of its sparse nature, while also being competitive in terms of mIoU, attaining almost perfect segmentations in most challenging scenarios.
@@
In Figure 2, we present our average performance for images arising from the three patterns in Figure 1. We vary $\ell$ and contaminate each image with Gaussian noise, $\sigma_{g}^{2}=\ell/100$, so they become increasingly noisier with size. These results demonstrate that our method scales well despite its spectral nature and generates high-quality segmentations on large images, even in challenging noise settings. The runtime growth depicted in Figure 2 reflects the linear complexity of our method.
In Figure 3, we evaluate the effect of $\mu$ on binary segmentation and qualitatively demonstrate our method’s performance and runtime on two sample images using their RGB intensities. Here, we note that $\mu$ has the effective role of ”blurring” the eigenvector and consequently smoothing the borders of the final segmentation, removing artifacts or boundary details as it increases. This clear and straightforward parameter interpretation is not present in the traditional formulations of the Normalized Cut problem [10], which entails that tuning it for a specific need is also simpler. Finally, we note our method’s low runtime in relatively large images without the need for any image approximation or interpolation despite its spectral nature.
(a) Image
(b) DFC
(c) DoubleDIP
(d) PiCIE
(e) SegSort
(f) Ours (RGB)
(g) Ours (DINO)
Fig. 4: Qualitative comparison of our method to other deep learning-based unsupervised segmentation methods on challenging BSDS500 images. Our proposed methodology is generally able to preserve borders well and, when using DINO features, qualitatively outperform other deep-learning based unsupervised segmentation techniques.
In Table 2, we quantitatively demonstrate the mIoU performance of our proposed method on BSDS images, comparing it to other unsupervised segmentation algorithms. We display results using RGB pixel intensities or DINO features interpolated to the pixel level. $\operatorname{\operatorname{NCut}}$ and W-Net perform surprisingly well compared to the other baselines and outperform many Deep Learning methods on BSDS. This can be seen as evidence of the power of spectral formulations to unsupervised segmentation tasks. Our method, when applied to RGB data, is only outperformed, slightly, by SegSort, which despite being unsupervised still makes use of pre-trained deep learning features. This means that our simple methodology when using just color data performs as well as other methods that rely on richer, but harder-to-train, features, demonstrating the suitability and appeal of our formulation to segmentation. On the other hand, when also making use of rich semantic features such as those from DINO, our method outperforms all depicted methods on BSDS.
Table 3 considers our segmentation methodology in the single object segmentation setting. Through this table, we note the overwhelming benefit of using spectral methods in the context of DINO features compared to other deep learning solutions. Furthermore, we again demonstrate the superior performance of our method on RGB data, achieving better performance results than many deep-learning solutions backed by rich semantic features. Indeed, it attains a state-of-the-art unsupervised segmentation performance on DUTS. Using DINO features on our algorithm usually improves its performance and also makes it achieve superior results on DUT-OMRON compared to DeepSpectral and TokenCut, our spectral DINO-backed comparison algorithms. We, however, fail to outperform these same methods on the ECSSD dataset, and future research will consider the reasons for this underperformance and approach different algorithmic improvements in our method to improve it.
Finally, Figure 4 qualitatively compares our method in its RGB and DINO formats against a few deep-learning-backed algorithms on challenging images from the BSDS dataset. Our method, especially when using DINO data, can detect low-contrast objects without losing border details. Again, this contrasts with the classical NCut results from [3], where much of segmentation borders are smoothed out due to their underlying graphical choice. Using extranodes, we qualitatively show that our method preserves borders well, while also having lower runtime and memory complexities.
5 Conclusion
@@
In this paper, we introduce a new spectral algorithm for image segmentation that overcomes some historical issues found in the $\operatorname{\operatorname{NCut}}$ framework and its related methods. We employ an efficient sparse graph, whose usage under the traditional $\operatorname{\operatorname{NCut}}$ optimization problem leads to an interpretable cut value and an efficient solver. We also show that our method, which is fully characterized by just one parameter, is fast and produces high-quality results under high-noise synthetic settings, and outperforms many state-of-the-art unsupervised image segmentation algorithms on popular image datasets, especially when backed by rich pixel-level features. These results demonstrate the effectiveness of using our sparse formulation for spectral image segmentation and provide further evidence for the successful use of spectral methods in self-supervised transformer data.